# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

### v4

@@ Abstract
We continue our analysis of Yablo’s coding of the liar paradox by infinite acyclic graphs, [Yab82].
The present notes are based on and continue the author’s previous results on the problem.
In particular, our approach is often more systematic than before.
@@ 1 Introduction
Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.
@@ 1.1 Outline
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. See also the author’s earlier versions on arXiv. In particular, the reader is referred there for a discussion of more complicated formulas attached to the nodes. We liberally use material from these sources.
All these papers are, of course, based on Yablo’s seminal [Yab82].
In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
In most cases, we will work with simple conjunctions (with and without negations).
In Section 2 (page 2), we discuss elementary contradictions (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction.
We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal. (See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].)
In both cases, we show that - within the chosen variant - the construction is necassary.
- (4)
There is an important property for the graphs, which we need, probably a sort of “richness”, see Remark 1.5 (page 1.5). This property holds trivially in Yablo’s construction by transitivity. In our approach, it will hold by the construction of a suitable injection, see Section 4 (page 4), Condition 4.1 (page 4.1) (2).
@@ 1.2 Basic Definitions and Facts
1.2.1 Terminology
Yablo’s construction
Yablo’s construction is an infinite directed graph with nodes and (negative) arrows for
Nodes stand for propositional variables, and the meaning of arrows is that variable occurs (negatively) in the formula attached to in most cases it expresses (We will identify for simplicity nodes with their attached variables.)
It is easy to see that no node may be true, nor false: Suppose is true, than all are false, so must be false, then contradiction, as all must be false. Suppose is false, then some must be true, again a contradiction (to the above).
Yablo Cell, Yablo Triangle
The basic contradiction structure has the form with the meaning (so and
We sometimes call the head, the knee, and the foot of the Yablo Cell (or Triangle).
Conditions (C1) and (C2)
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 3.1 (page 3.1), (3).)
1.2.2 The Logical Side
On the logics side, we work with propositional formulas, which may, however, be infinite.
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 1.6 (page 1.6).
@@ 1.2.2 The Logical Side
When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
We will work here with disjunctive normal forms, i.e. with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
Fact 1.1
Let where the are propositional variables or negations thereof.
@@
Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
Fact 1.2
Let and we want to show that is contradictory, then it suffices to show that a subset of the is inconsistent - this corresponds to the fact that (X any truth value.)
Let and we want to show that is contradictory, then we have to show that ALL are contradictory. Again, this corresponds to
The following example illustrates the propagation of required contradictions.
Example 1.1
Consider
As has to be contradictory, both components and have to be contradictory.
As has to be contradictory, all components, etc., have to be contradictory. For this, it is sufficient and necessary that in each component, one element is contradictory. This is equivalent to the condition that (both and or (both and are contradictory.
Thus
For both
and
have to be contradictory, this is done as usual by
and
For we have to attach a contradiction at either
(-1) and
or
(-2) and
We thus have to examine the combinations
(-1)
(-2)
(-1)
(-2)
The combinations (-1)) and (-2)) are as usual, the other ones are new.
It seems that there might be a problem.
In the cases and seems to be independent from and both solvable by classical logic, a contradiction. But all cases are parts of and and (-1)) and (-2)) need simultaneous treatment of the and the - part.
So the confusion is only transient.
(In the case we have no independent parts, so the “problem” just discussed will not occur.)
1.2.3 The Graph Side
We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
Definition 1.1
Nodes stand for propositional variables.
If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for more precisely.
Example 1.2
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.2.1 (page 1.2.1), (2)):
Let so so we have consistency, and it does not work for the construction.
Example 1.3
Consider the basic construction of a contradiction (used by Yablo and here, defined “officially” in Definition 2.1 (page 2.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
@@ 1.2.3 The Graph Side
If all arrows are negative, then iff the length of is even.
Definition 1.3
We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
Remark 1.3
Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
We saw already in Example 1.3 (page 1.3) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
Fact 1.4
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
(Trivial, we have only 2 truth values).
Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ). Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
Above properties generalize to any loops of odd length (with more paths).
This does not hold when the paths may branch again after meeting, as the next Example shows.
Example 1.4
(Example 7.4.2 in [Sch22].)
Let
then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 1.5
Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
Thus For any we have a contradiction by and for any and for any for a suitable
Example 1.6
This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “procrastination branch” or “escape branch”. This modification of the Yablo structure has one acceptable valuation for
Let as usual, and introduce new
Let with
If then by so, generally,
if then and
If then so if etc., so, generally,
if then or
But is possible, by setting and for all
Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “procrastinating” branch.
Diagram 1.1
Escape Path
Remark 1.5
For a representation result, we need enough paths in the graph considered to form contradictions. This is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition directly. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
1.2.4 Interplay of the Graph and the Logical Side
We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
We need infinite depth and width in our constructions:
Fact 1.6
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
the construction needs infinite width, i.e. the logic cannot be classical.
Proof
Let be a minimal element, then we can chose an arbitrary truth value and propagate this value upwards. If there are no infinite descending chains, we can do this for the whole construction.
The logic as used in Yablo’s construction is not compact:
Trivial. (Take This is obviously inconsistent, but no finite subset is.).
It is impossible to construct a Yablo-like structure with classical logic:
Take an acyclic graph, and interpret it as in Yablo’s construction. Wlog., we may assume the graph is connected. Suppose it shows that cannot be given a truth value. Then the set of formulas showing this does not have a model, so it is inconsistent. If the formulas were classical, it would have a finite, inconsistent subset, Define the depth of a formula as the shortest path from to this formula. There is a (finite) such that all formulas in have depth Give all formulas of depth (arbitrary) truth values, and work upwards using truth functions. As the graph is acyclic, this is possible. Finally, has a truth value.
Thus, we need the infinite
The following example illustrates (partially) unfinished contradictions, and resulting escape possibilities.
Example 1.7
There is a fundamental difference between the contradictions (for and
Consider the case The path is blocked at So we cannot go to by two contradicting possibilities. The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility. So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative. This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape. So a negative start at the original diagram leads to a positive start in the appended diagram.
Diagram 1.2
Diagram Postponed Contradictions
1.2.5 A Third Truth Value
See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
Remark 1.7
We consider a third truth value, in addition to and
means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
@@ 1.3 Strategy
The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances.
1.3.1 Expressing “x is false” In an Acyclic Graph
In an acyclic graph, we cannot express something like we have to use different means to express that is impossible.
As we can always fill in truth values (including from the bottom, we cannot force by looking upwards of (or sideways, i.e. there is no monotone path connecting with some So we have to look downward.
We may use or if we have the constant We will see later, that this not interesting, see Definition 2.1 (page 2.1) and Remark 2.1 (page 2.1). This is a contradiction made of (graphically speaking) two paths of length one each. Interesting contradictions are made of slightly more complicated paths, like the Yablo Cell meaning
Basically, the only way to express that is impossible, is by the logical formula
A slightly different argument: How can we express that is wrong in a finite acyclic graph? Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
a disjunction of conjunctions of elements in , or their negations. By acyclicity,
This argument is not totally satisfactory. For this reason, we reformulate:
We examine graphs, where there existe a node such that as well as -x will lead to a contradiction.
@@ 1.3.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
So we want to express by an infinite graph that both and are inconsistent. As we saw in Section 1.3.1 (page 1.3.1), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
@@
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3).
A summary: we have to look at instances.
@@
1.4 The “Right” Level of Abstraction
1.4.1 Introduction
The basic elements in Yablo’s construction are negative arrows, from which Yablo Cells are built. We show here that we can build arbitrarily complex structures equivalent to negative arrows, and as a matter of fact to any propositional logical operator. This suggests that the right level of abstraction to consider more general Yablo-like structures is not the level of single arrows, but rather of paths.
Remark 1.8
In the diagrams in Diagram 1.3 (page 1.3) the inner path is barred by , and the outer path contains 3 or 2 negations.
@@ 1.4.1 Introduction
We define negations and their type.
A negation (diagram) is a diagram all of whose arrows are of the kind.
We define negation diagrams, or, simply, negations, and their type.
An arrow is a negation of type 0.
all negations inside are of type and at least one negation inside is of type
Remark 1.9
Thus, it seems very difficult to describe Yablo type diagrams on the level of single arrows. It seems to be the wrong level of abstraction.
We now give some examples.
1.4.2 Some Examples
Example 1.8
Negation (1)
so and
Negation (2)
so and continue as above.
Identity
so and
TRUE
Thus, and
so and
Diagram 1.3
Propositional Formulas 1
Diagram 1.4
Propositional Formulas 2
1.5 Prerequisites
We work with pure conjunctions of positive or negative propositional variables, like except for the following:
we may use a constant TRUE or
we may use formulas of the type
so the following are also permitted: etc.
This is used only to “neutralize” unwanted parts of the graph we characterize, see Fact 4.4 (page 4.4).
- (2)
- (3)
- (4)
@@ 2.1 Introduction
Elementary cells have to be simple constructions providing contradictions (in some cases, e.g. when the origin, usually called or so, cannot have a certain truth value, as this will lead, within the cell, to a contradiction).
Thus
They must be contradictory, and thus exclude at least one instance.
In addition, they should be simple, e.g. they will not contain a sequence or so without branching, as we could replace this with
(We will see later how to relinquish this requirement.)
More importantly, we must be able to construct a Yablo-like structure, built up by this type of elementary cell only (and perhaps some simpler auxiliary structures). They must serve as building blocks for a suitable stucture. This prevents “cheating”, letting other constructions do the crucial work.
- (4)
@@ 2.2 Basic Discussion
See Section 7.3 in [Sch22] for more material.
We try to describe here the basic constructions of contradictions.
As said in Fact 1.1 (page 1.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Section 2.1 (page 2.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
@@
An escape problem with for the trivial contradiction, so (See Definition 2.1 (page 2.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
@@
There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
@@
Example 2.1
We consider now some simple, contradictory cells. They should not only be contradictory for the case but also be a potential start for the case without using more complex cells. (Otherwise, we postpone the solution, and may forget the overly simple start.)
We will consider in Section 2.5 (page 2.5) a more complicated contradiction cell (“Diamond”), and discuss here only simple cells.
@@
The cell with 2 arrows.
It corresponds to the formula graphically, it has a positive and a negative arrow from to so exactly one of and is negative.
If is positive, we have a contradiction.
Of course, appending at a Yablo Cell (see below, case (2.3)) may be the beginning of a contradictory structure, but this is “cheating”, we use a more complex cell.
Cells with 3 arrows.
In the graph with not annotated arrows, they have the form -z, -z, where the lines may stand for or
@@
This will be discussed in detail in the rest of the paper. But we see already that both paths, and change sign, so and will be positive, appending the same type of cell at and solves the problem (locally), and offers no escape.
We call sometimes the head of the Yablo cell, its foot, and its knee.
If we combine Yablo Cells, the knee for one cell may become the head for another Cell, etc.
@@
Upper part:
Lower part:
@@
Here,
Diagram 2.1
Simple Contradictions
Diagram 2.2
Diagram ESC-1
Diagram 2.3
Diagram ESC-2
Diagram 2.4
Diagram ESC-3

### v5

@@ Abstract
We show that the existence of a node with attached formula in an acyclic graph such that neither nor its negation are consistent is equivalent to the existence of subgraph equivalent to Yablo’s construction.
@@ 1.1 Outline
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. See also [Sch25]. (Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.) We liberally use material from these sources. All these papers are, of course, based on Yablo’s seminal [Yab82].
Prerequisites:
All graphs considered will by acyclic.
- (1.2)
For ease of exposition, we introduce a third truth value, as abbreviation for “ and are inconsistent”.
We will use only one type of contradictory cell in a graph. This prevents “cheating”.
Contradictions in a graph will begin immediately, and not after some arbitrary path. This simplifies constructions, and is not an important restriction.
- (1.6)
In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent. In Section 1.3 (page 1.3), we discuss our strategy, and discuss the antagonistic conditions (C1) and (C2) and how to satisfy them in an inductive construction. (We need infinite depth and width in the construction, see e.g. [Sch25], Fact 2.4, or [Sch23b], Fact 1.4).
In Section 2 (page 2), we discuss elementary contradictions (cells). Essentially, these are variations of Yablo’s triangles. The trivial contradiction, see Definition 2.1 (page 2.1), is too simple, it has an “escape possibility”, i.e., combining them will leave one possibility without contradictions. See Definition 1.6 (page 1.6). Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework. See Section 2.5 (page 2.5). We also show that all sides of the triangles have to be negative, see Remark 2.2 (page 2.2). Section 2.3 (page 2.3) discusses further details of these triangles. We generalize to triangles formed by paths in Section 2.4 (page 2.4).
In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction. We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal. This is perhaps the central result of the paper. This fact is essentially re-used in Section 4 (page 4). (See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].) In both cases, we show that the construction is necessary.
In Section 4 (page 4), we show that
a graph without valuation can be given a valuation such that there is an element such that neither nor are consistent, if there is a suitable injection from Yablo’s structure into and a partial valuation for the nodes of the paths of the injection, such that contradictions are preserved. We complete the partial injection by a trivial extension.
conversely, if there is a graph with (complete) valuation with some node such that neither nor are consistent, then we can construct a subgraph which corresponds again to an injection of the Yablo structure. We use here again the minimality of Yablo’s construction, this time imside the given graph.
@@ 1.2 Basic Definitions and Facts
We consider directed, acyclic graphs. They will usually have one root, often called Nodes and arrows may be labelled.
If the nodes are labelled, we identify nodes with their labels, unless said otherwise. The labels of nodes stand for propositional variables or constants The propositional variables will be unique in a given graph. Moreoever, to is associated a formula the direct successors of a node determine the propositional variables or constants occurring in The theory of a graph is the union of its a node in the graph. We sometimes write for the theory of a graph
We sometimes write if occurs negatively in
In many cases is a conjunction of positive or negative occurrences of the direct successors of If not, we will say so expressly. For instance, the graph will express that we often write instead of
Definition 1.1
The pure graph, i.e. just nodes and arrows, tells only which variables occur in the associated formulas We call the annotations or valuations, the pure graph graph without valuations or annotations, and the full graph annotated graph or graph with valuations, etc. If is the pure graph, will sometimes be an annotated version of and conversely.
Definition 1.2
Let be a node in a fixed graph as above.
We say that is contradictory (relative to iff the assumption leads to a contradiction in Likewise for is contradictory, iff the assumption is contradictory in this theory.
Truth values:
We set iff is contradictory, and iff not, likewise for
We introduce a third truth value which is an abbreviation for and - relative to of course. See Remark 1.2 (page 1.2).
When the context is clear, we sometimes simplify to etc.
Contradictions via paths - we will change between the graph and the logical formulation, whatever seems intuitively better:
Such contradictions have in a graph the following form: there is a node and a node and two paths from to which branch at and join again at and they contradict each other, i.e. one branch results in the other in
These are the basic contradictions.
- (3.2)
Consider again Diagram 3.1 (page 3.1), part (2.1). The whole diagram is not contradictory, as we opened a new possibility at which becomes impossible only in part (3) of this diagram. So, as a detail, the paths are contradictory, but the possibility is closed only in (3), so the contradiction of the whole diagram is only complete in (3). If you prefer to think probabilistically, in part (2.1) of the cases are contradictory etc. See also Section 1.3.1 (page 1.3.1) for opening and closing new possibilities.
Definition 1.3
(Yablo, see [Yab82].)
We call the following graph the Yablo structure, YS:
Written as graph: We use negated arrows, and positive arrows, in graphs as this improves readability considerably (see Notation 1.1 (page 1.1).)
Written as set of propositional formulas:
Contradictions in YS:
Suppose is a node in the YS. Then and which is contradictory.
But starting with will also result in a contradiction:
suppose is true, then is contradictory, as we just saw, so is contradictory, too.
Note that we abused notation here, where may be an arrow in a graph, or an implication in a logical formula. Context will tell.
The contradictions have the form this is the basic contradiction “cell”, we call such cells Yablo cells, ’s, or Yablo triangles.
We sometimes call the head, the knee, the foot of such ’s. This is relative to the cell discussed. When we combine such cells, the knee or foot of one cell may become the head of the next cell, etc.
Remark 1.1
Let where the are propositional variables or negations thereof, thus, is in disjunctive normal form, DNF.
@@
Example:
Consider the Yablo cell with each formula associated with of the type We then have 16 possible combinations like etc., 7 are in 9 in is in the 9 combinations of and the 9 combinations of By is in the 7 positive combinations of so we have again a contradiction, as usual.
(Alternatively, we may simplify, and work instead of with many combinations, e.g. just with the totally positive combination for and with the totally negative combination
The generalization to arbitrary DNF’s is trivial.
Thus, working with DNF’s changes nothing compared to the Yablo construction, working with All we need is a suitable negation.
Remark 1.2
The truth value has the following properties:
(and, of course
the following order is compatibel with and i.e. for any truth values and
This completely describes the behaviour of
This is the only 3-valued logic with the property and compatibility with and It is thus identical to the Strong Kleene logic Ks, as was pointed out by a referee.
Proof:
For more details, see [Sch25]
trivial
Consider two copies of ’s, YS and YS’, beginning with and respectively, and the following two structures (a) with (b) with
Consider a new origin, and arrows from to as follows: and and
Details are left to the reader.
Easy.
@@ 1.3 Strategy
Definition 1.5
Conditions (C1) and (C2)
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction.
@@ 1.3.1 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
So we want to express by an infinite graph that both and are inconsistent. See Definition 1.2 (page 1.2). Therefore, we treat such contradictions as basic building blocks, as Yablo did.
@@
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3), we have to look at instances.
@@
Example 1.1
This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “escape branch” or path. See below, Definition 1.6 (page 1.6). This modification of the Yablo structure has one acceptable valuation for
Let as usual, and introduce new
Let with
If then by so, generally,
if then and
If then so if etc., so, generally,
if then or
But is possible, by setting and for all
Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “escape” branch, see below.
@@
When we try to make a structure contradictory for some value of the origin, but find a way (a valuation) to avoid or escape a contradiction, we call such a valuation an escape path, branch, or possibilty. This is then a model consistent with the theory of the graph.
This seems an intuitive description how the structure fails to be contradictory for this valuation.
@@ 2.1 Introduction
We try to generalize Yablo cells to some other types of cells.
(C1) They will have the form of two paths which branch say at and meet again at (but not before), and are contradictory.
(C2) Both branches will have to lead to new contradiction cells if
We require that it is possible to construct a structure with such cells starting, say at such that and are contradictory.
In addition, we require that this is possible using only the contradiction cells of this type. This prevents “cheating”, where the real work is done by some other type of cell.
We will impose syntactical limitations on the formulas attached to the nodes in the cells.
@@ 2.2 Basic Discussion
- (1)
@@
There are, whatever the inner structure, just four possibilities: it might be equivalent to
The first two cases are trivial, as we have the possibilities or so is always a possibility.
Suppose Then
Suppose Then
@@
If we continue with a copy of now grafting on etc., we may not achieve a desired result (whereas a different may achieve it).
Remark 2.2
We consider here cells with three arrows.
@@
In the graph with not annotated arrows, they have the form , , where the lines may stand for or
@@
This will be discussed in detail in the rest of the paper. But we see already that both paths, and change sign, so and will be positive when is negative, appending the same type of cell at and solves the problem (locally), and offers no escape.
@@
Consider the theory indicates an escape path.
Consider the theory indicates an escape path.
@@
Consider the theory indicates an escape path.
See also the discussion of “Saw Blades” in [Sch22], Section 3.
2.3 Values for y and z in Yablo Cells
We consider here what happens when we give and truth values In addition, we consider the variants and